OpenStudy (watermelon14):

a rope is swinging in such a way that the length of the arc is decreasing geometrically. If the first arc is 27 feet long and the fourth arc is 8 feet long, what is the length of the second arc?

OpenStudy (watermelon14):

@jim_thompson5910

OpenStudy (watermelon14):

my equation i came up with this 27(2/4)^n-1 is that not right?

jimthompson5910 (jim_thompson5910):

`length of the arc is decreasing geometrically` so if 27 is the first arc length then 27*r is the second arc length where r is some value in the interval 0 < r < 1 the third arc is 27*r*r = 27*r^2 the fourth arc is 27*r^2*r = 27*r^3, and we're given it's equal to 8, so 27r^3 = 8 making r = ???

OpenStudy (watermelon14):

is it .6666

jimthompson5910 (jim_thompson5910):

or 2/3

jimthompson5910 (jim_thompson5910):

so 27*r = 27*(2/3) = ???

OpenStudy (watermelon14):

12

jimthompson5910 (jim_thompson5910):

no

OpenStudy (watermelon14):

oh i did it squared to find the second

OpenStudy (watermelon14):

18

jimthompson5910 (jim_thompson5910):

yep the second arc is 18 ft

OpenStudy (watermelon14):

wait wouldn't it be 12 because it is squared for 2nd

jimthompson5910 (jim_thompson5910):

Common ratio: r = 2/3 first arc = 27 ft .... given second arc = (first arc)*(r) second arc = (27)*(2/3) second arc = 18 ft third arc = (second arc)*(r) third arc = (18)*(2/3) third arc = 12 fourth arc = (third arc)*(r) fourth arc = (12)*(2/3) fourth arc = 8 ... also given and this confirms what we're given

OpenStudy (watermelon14):

oh i see but idk why 18 isn't one of my answer choices

jimthompson5910 (jim_thompson5910):

there's a typo somewhere I'm assuming

OpenStudy (watermelon14):

oh maybe. but u seem correct

jimthompson5910 (jim_thompson5910):

The only thing you can do is bring it up with the teacher

OpenStudy (watermelon14):

yeah i'll ask. thank you!

jimthompson5910 (jim_thompson5910):

no problem